OpenStudy (anonymous):

Write (1+3i)(5-i) in standard form ? :)

hartnn (hartnn):

can you distribute ? like (a+b)(c+d) = ac+ad+bc+bd

OpenStudy (anonymous):

8+ 14i ? (:

hartnn (hartnn):

let me check, 5+15i-i-3i^2 = 5+14i+3 = 8+14i wow! absolutely correct :D

OpenStudy (anonymous):

yay! :D